OpenStudy (funnyguy2):

What is the ratio of x to y? 2/3 = x/3x + y they are fractions

OpenStudy (funnyguy2):

a. -3/2 b. -2/3 c. 2/3 d. 3/2

OpenStudy (campbell_st):

start by cross multiplying \[2(3x + y) = 3x\] simplify and then collect like terms

OpenStudy (funnyguy2):

2(x + y) = x

OpenStudy (campbell_st):

\[\frac{2}{3} = \frac{x}{3x + y}\] if you equate the numerators x = 2 substitute it into the denominator and 3 = (3x + y) find y then you get the ratio

OpenStudy (funnyguy2):

d?

OpenStudy (campbell_st):

the ratio is x in the numerator and y in the denominator x = 2 is the top of the fraction... so what was y, when you calculated it?

OpenStudy (funnyguy2):

-3

OpenStudy (campbell_st):

great so what is the ratio as a fraction... x on top y on the bottom